OpenStudy (anonymous):

The roots of the quadratic equation 8x^2-4kx+27=0 are a and a^2 Show that k=7.5 and a=3/2 Is it possible to use the discriminant of the equation to find the value of k?

OpenStudy (inkyvoyd):

Hey, don't close and reopen a question just to bump it. It is against the rules, and not fair to those who bump their questions instead of closing and opening. It is also cheating the system because you are getting extra credit for asking the same question.

OpenStudy (inkyvoyd):

If your question is not answered, you may want to bump it until it is, and tag a few people, asking nicely for help.

OpenStudy (shubhamsrg):

use this: in ax^2 +bx +c=0 sum of roots = -b/a and product of roots = c/a

OpenStudy (anonymous):

ya. use Vieta theorem for ax^2+bx+c=0 sum of roots = -b/a and product of roots = c/a so product of roots a*a^2=27/8, a=3/2 and substitude a=3/2 as a root into this equation u can have k=7.5